if numerator of fraction is increased by 1 and denominator is decreased by 1, then fraction becomes 1/2. if numerator is increased by 1 less than twice of itself while denominator is increased by 2 more than thrice of itself, then fraction becomes 1/6. the original fraqtion is ?

Grade:10

2 Answers

Vijay Mukati
8 years ago
Assume orignial fraction to be x/y.

According to 1stcondition, (x+1)/(y-1) = ½
According to 2ndcondition, (x + (2x-1))/(y+(3y+2)) = 1/6

Onsolving above two conditions, you will get x = 2 and y = 7. therfore, the franction is 2/7

Thanks.
Palmakir
8 years ago
Let the fraction be x/y
(x+1)/(y-1) = ½
(x+2x-1)/(y+3y+2) = 1/6
Solve this two equations......
x=2,y=7
Therefore the fraction is 2/7